is the following trinomial a perfect square?
x^2-12x+36

Answer by ewatrrr(24785)

 
Hi,
is the following trinomial a perfect square? YES
x^2-12x+36 = (x-6)(x-6)= (x-6)^2
F First terms x^2
O Outside terms -6x
I Inside terms -6x
L Last terms 36